OpenStudy (anonymous):

Which of the following sets of numbers is a Pythagorean triple? 4, 5, 6 8, 15, 17 5, 12, 14 8, 24, 25

OpenStudy (ivettef365):

use formula to see which one agrees a^2 + b^2 = c^2

OpenStudy (ivettef365):

lets check the first 4, 5, 6 4^2 + 5^2 = 6^2 16 + 25 = 36 41 = 36 NOT GOOD, keep checking

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Not applicable to this problem, but for future similar problems: 3, 4, 5 is the best know pythagorean triple and any multiple of that. So... 3,4,5 \(\rightarrow 3^2+4^2=5^2 \rightarrow 9+16=25 \rightarrow 25=25\) 6,8,10\(\rightarrow 6^2 + 8^2= 10^2 \rightarrow 36 + 64 = 100 \rightarrow 100=100\) 9,12,15\(\rightarrow 9^2+12^2=15^2 \rightarrow 81+144 = 225\rightarrow 225=225\) And the pattern continues to infinity. Checking for these first may save you time on a test.